Speed Racer accelerates from rest at a constant rate of 6.5 m/s2. How fast will the racer be going at the end of 8 s?


1
Expert's answer
2022-05-05T13:56:42-0400
"v=v_0+at\\
=0+6.5*8=52\;\rm m/s"